OpenStudy (magbak):

MAGDY AND @Jamierox4ev3r TRAINING GROUNDS DO WHAT EVER YOU WANT TO DO RIGHT HERE WHEN EVER YOU WANT AND WHAT EVER YOU WANT. (NO INAPPROPRIATE LANGUAGE PLEASE)

HanAkoSolo (jamierox4ev3r):

OMG :P

OpenStudy (magbak):

OUR TRAINING GROUND FOR LATEX

HanAkoSolo (jamierox4ev3r):

idk what jus happened but okay \(\huge\mathbb{L}\mathbb{E}\mathbb{G}\mathbb{I}\mathbb{T}\mathbb{!}\)

OpenStudy (magbak):

\(\LaTeX\)

OpenStudy (magbak):

\(\LEGITE\)

OpenStudy (magbak):

Yea, LA TE X is Fun Yea, LA TE X is Fun

OpenStudy (magbak):

@Jamierox4ev3r

OpenStudy (e.mccormick):

\(\TeX{T}\) Messages in funky type!

OpenStudy (abb0t):

No.

OpenStudy (swag):

lol

OpenStudy (ybarrap):

$$ \begin{array}{cc} \huge \ddot \smile\\ \bowtie&\\ \scriptsize^\circ\!&\!\!\!\!\!\!\!\!\!\!\small ^\mathbb{LEGIT!}\\ \scriptsize^\circ\!&\!\\ \end{array} $$

OpenStudy (e.mccormick):

\[ \,\Huge{\overset{\frown}{\normalsize \left( \begin{matrix} \hat\odot \; \hat\odot\\ \Downarrow\\ \Large \color{#CD5555}{\smile} \end{matrix} \right)}}\normalsize \\ \begin{array}{lcr} /&\bowtie&\backslash\\ |&\scriptsize^\circ&|\\ |&\scriptsize^\circ&|\\ \end{array} \]

OpenStudy (anonymous):

{huge/red/ what are you doing?/huge}

OpenStudy (anonymous):

what the deuce? Sorry, I really am bad at LaTeX

OpenStudy (e.mccormick):

`\( LaTeX code goes here \)` or `\[ LaTeX code goes here \]` The `\(\) \[\]` parts are critical.

OpenStudy (e.mccormick):

And the code slashes are \ `\(\huge \color{red}{\text{Hi there!}}\)` makes: \(\huge \color{red}{\text{Hi there!}}\)

OpenStudy (anonymous):

\(huge\color{green}{\text{ Hello my friend!}\)

OpenStudy (anonymous):

much confuzzled complicated much no understand

OpenStudy (e.mccormick):

You missed the \ in front of \huge

OpenStudy (anonymous):

Oh gosh. Thanks! \huge\color{green}{\text{Hello my friend!}

OpenStudy (anonymous):

I quit LaTeX. Alright. Bye OS. =I

OpenStudy (e.mccormick):

\(\huge\color{green}{\text{Hello my friend!}}\) `\(\huge\color{green}{\text{Hello my friend!}}\)` Ot can take getting used to.

thomaster (thomaster):

\(\Huge\bf\color{#FF0000}\thicksim \LARGE \color{#FF9200}S\color{#FFDB00}i\color{#FFff00}x\ \color{#B6ff00}M\color{#6Dff00}o\color{#00ff00}n\color{#00ff49}t\color{#00ffDB}h\color{#00ffff}s\ \color{#00DBff}L\color{#0092ff}a\color{#0049ff}t\color{#0000ff}e\color{#2400ff}r\ \color{#6D00ff}x\color{#B600ff}D\)

OpenStudy (e.mccormick):

Well, it is an open training ground....

OpenStudy (ybarrap):

$$ \Huge \mathfrak{ \hspace{100pt}\bbox[5pt,#ffffff,border:10px solid #0a0a0a]{\sqrt{-1}\cdot2^3\cdot\Sigma\cdot\pi} } $$ \(\hspace{150pt}\text{That is}\) $$ \hspace{125pt}\bbox[5pt,#ffffff,border:10px solid #bbbbbb]{\text{I Ate Some Pie}} $$ $$ \Huge{ \text{And Loved IT!}} $$

OpenStudy (barrelracer011):

\(\huge \color{purple}{\text{Hi there!}}\)

OpenStudy (anonymous):

/(/huge/color{purple}{/text{BLUB}}/)

OpenStudy (ybarrap):

Change `/(/huge/color{purple}{/text{BLUB}}/)` to `\(\huge\color{purple}{\text{BLUB}}\)` I would do it but I want you to get the excitement from seeing it yourself

TheSmartOne (thesmartone):

LOL that pun though @ybarrap xDD